OpenStudy (andriod09):

I need help on a Factoring problem. Its factoring a fraction, on top of another factoring fraction. Its in the comments

OpenStudy (andriod09):

\[\huge\frac{\frac{30x^2+27x+6}{9x^2-49}}{\frac{30x+15}{30x-70}}\]

OpenStudy (anonymous):

The size of those fraction bars is important. The top two appear to be the same length, and the third is smaller. Correct?

OpenStudy (strawberry17):

To factor the numerator of the top fraction, you would divide each number by 3 and get: 3(10x^2 + 9x +2)

OpenStudy (andriod09):

no, the fracton bar on the top is supposed to be in the middle, i.e. it should be the top fraction normal bar, then solid black bar, then bottom fraction normal. @strawberry17 how would i do that?

OpenStudy (anonymous):

\(\Huge \frac{30x^2+27x+6}{9x^2-49} \div\frac{30x+15}{30x-70}\) \(\Huge =\frac{30x^2+27x+6}{9x^2-49} *\frac{30x-70}{30x+15}\)

OpenStudy (andriod09):

its supposed to be: \[\frac{30x^2+27x+6}{9x^2−49}\] over with the solid black line, \[\frac{30x+15}{30x−70}\]

OpenStudy (anonymous):

Like I put, Andriod?

OpenStudy (andriod09):

i don't understand what you did.

OpenStudy (anonymous):

One fraction divided by another fraction.

OpenStudy (andriod09):

i don't know. i tried to put it like it is in the book. thats what it says,

OpenStudy (andriod09):

yes, but i get very confused with fraction.

OpenStudy (anonymous):

Okay, Andriod, when we divide by a fraction, it's the same as multiplying by the reciprocal. That is, \(\div 2\) is the same as \(\large *\frac{1}{2}\) \(\large \div\frac{2}{3}\) is the same as \(\large *\frac{3}{2}\)

OpenStudy (anonymous):

etc. So when we see \(\huge\div \frac{30x+15}{30x-70}\) it's easier for us if we rewrite it as \(\huge *\frac{30x-70}{30x+15}\)

OpenStudy (andriod09):

ik that much, but i don't know what to do wit hthe FRACTIONS, i.e. i know the KFC method, but what/how do i factor the fraction?

OpenStudy (strawberry17):

@andriod09 do you understand how I factored out 3 from the numerator of the top fraction?

OpenStudy (anonymous):

What does KFC stand for?

OpenStudy (andriod09):

K=keep F=flip C=change @strawberry17 i have not a clue.

OpenStudy (anonymous):

Andriod, look for a common factor. With \(30x^2 + 27x + 6\) you can notice that 30=3*10 27= 3*9 6= 3*2 so they have a common factor of 3.

OpenStudy (anonymous):

I agree, Strawberry. Rewriting it is good, but it's good to keep it factored as it is.

OpenStudy (andriod09):

i don't understand. \[:(\]\[:L\]\[:/\]

OpenStudy (andriod09):

yes. i undertand what factoring is. I don't konw how to factor a fraction. hence, why this post is here.

OpenStudy (andriod09):

FACTORING A FRACTION IS ALL I DO NOT KNOW. I KNOW HOW TO FACTOR SOMETHING LIKE THIS: \[x^2+3-27\]

OpenStudy (anonymous):

Factoring a fraction is not different from factoring normally. Just look at the top and factor it normally. Then look at the bottom and factor it normally.

OpenStudy (andriod09):

I DON'T KNOW HOW. :/

OpenStudy (andriod09):

I DON'T KNOW HOW TO FACTOR A FRACTION.

OpenStudy (andriod09):

just tell me how to do this problem please, i am getting really annoyed.